# Prophet: Prompting Large Language Models with Complementary Answer Heuristics for Knowledge-based Visual Question Answering

### v3

@@ Abstract
Knowledge-based visual question answering (VQA) requires external knowledge beyond the image to answer the question. Early studies retrieve required knowledge from explicit knowledge bases (KBs), which often introduces irrelevant information to the question, hence restricting the performance of their models. Recent works have resorted to using a powerful large language model (LLM) as an implicit knowledge engine to acquire the necessary knowledge for answering. Despite the encouraging results achieved by these methods, we argue that they have not fully activated the capacity of the blind LLM as the provided textual input is insufficient to depict the required visual information to answer the question. In this paper, we present Prophet—a conceptually simple, flexible, and general framework designed to prompt LLM with answer heuristics for knowledge-based VQA. Specifically, we first train a vanilla VQA model on a specific knowledge-based VQA dataset without external knowledge. After that, we extract two types of complementary answer heuristics from the VQA model: answer candidates and answer-aware examples. Finally, the two types of answer heuristics are jointly encoded into a formatted prompt to facilitate the LLM’s understanding of both the image and question, thus generating a more accurate answer. By incorporating the state-of-the-art LLM GPT-3 [1], Prophet significantly outperforms existing state-of-the-art methods on four challenging knowledge-based VQA datasets. To demonstrate the generality of our approach, we instantiate Prophet with the combinations of different VQA models (i.e., both discriminative and generative ones) and different LLMs (i.e., both commercial and open-source ones).
@@
Visual Question Answering (VQA), large language models (LLMs), knowledge-based VQA, multimodal learning.
@@ 1 Introduction
Recent advances in deep learning have enabled substantial progress in visual question answering (VQA) which requires a machine to answer free-form questions by reasoning about given images. Benefiting from large-scale vision-language pretraining, the state-of-the-art methods have even surpassed human level on several representative benchmarks [2, 3, 4]. Despite the success of these methods, their reasoning abilities are far from satisfactory, especially when external knowledge is required to answer the questions. In this situation, the task of knowledge-based VQA is introduced to validate models’ abilities to leverage external knowledge. Early knowledge-based VQA benchmarks additionally provide structured knowledge bases (KBs) and annotate required knowledge facts for all the questions [5, 6]. More recently, benchmarks emphasizing on open-domain knowledge have been established [7, 8], which means KBs are no longer provided and any external knowledge resource can be used for answering. We focus on the task with open-domain knowledge in this paper.
A straightforward solution for knowledge-based VQA is to retrieve knowledge entries from explicit KBs, e.g., Wikipedia and ConceptNet [9]. Then, a KB-augmented VQA model performs joint reasoning over the retrieved knowledge, image, and question to predict the answer [10, 11, 12, 13, 14]. However, the performance of these retrieval-based approaches is limited for two reasons: (i) the required knowledge may not be successfully retrieved from the KBs; and (ii) even if the required knowledge is retrieved, plenty of irrelevant knowledge is inevitably introduced, which hampers the learning of VQA models.
@@
Figure 1: Conceptual comparisons of three knowledge-based VQA frameworks using a frozen LLM model, e.g., GPT-3 [1]. While PICa [15], KAT [16], and REVIVE [17] directly feed the caption (C) and question (Q) into the LLM as the prompt, we argue that the information they provide for the LLM is insufficient thus cannot fully activate the LLM’s potential. In contrast, our Prophet learns a vanilla VQA model without external knowledge to produce answer heuristics, which endows the LLM with richer and more task-specific information for answer prediction. In contrast to the counterparts that resort to specific VQA models and LLMs, our Prophet is more general that can be instantiated with the combinations of different VQA models (i.e., discriminative [18] and generative ones [19]) and different LLMs (i.e., commercial [1] and open-source ones [20, 21]).
@@
    The generated captions cannot cover all the necessary information in the image. Consider the example in Fig. 1, the caption “a group of people walk in a city square” contributes nothing to answering the question “what fruit comes from these trees”. In this situation, the LLM has to make an aimless and biased guess to answer the question.
@@
    LLMs like GPT-3 employ a few-shot learning paradigm that requires a few in-context examples to adapt to new tasks. Therefore, the choice of these examples is critical to model performance. As reported in [15], all its example selection strategies achieve far inferior performance to the oracle strategy that uses the similarity of ground-truth answers.
@@
In this paper, we present Prophet—a conceptually simple yet effective framework designed to prompt LLMs with answer heuristics for knowledge-based VQA. By answer heuristics, we mean some promising answers that are presented in a proper manner in the prompt. Specifically, we introduce two types of complementary answer heuristics, namely answer candidates and answer-aware examples, to overcome the limitations in (i) and (ii), respectively. Given a testing input consisting of an image and a question, the answer candidates refer to a list of promising answers to the testing input, where each answer is associated with a confidence score. The answer-aware examples refer to a list of in-context examples, where each example has a similar answer to the testing input. Interestingly, these two types of answer heuristics can be simultaneously obtained from any vanilla VQA model trained on a specific knowledge-based VQA dataset. A schematic of Prophet is illustrated at the bottom of Fig. 1.
@@
A preliminary version of this manuscript was published in [22]. Based on that version, we have made the following contributions to further improve the performance and validate the generality of Prophet: (i) we investigate diverse types of VQA models, including the classical discriminative models trained from scratch and the latest generative VQA models pretrained on large-scale corpus; (ii) we expand the used LLM from the commercial GPT-3 model to a wide range of open-source models; (iii) apart from OK-VQA and A-OKVQA, we conduct more experiments on two other knowledge-based VQA datasets, namely ScienceQA [23] and TextVQA [24]. The source code is made available here [^1] https://github.com/MILVLG/prophet. We hope these studies may serve as a new baseline to inspire future research on knowledge-based VQA and universal vision-language learning.
@@ 2 Related Work
Visual Question Answering (VQA). VQA has been of growing interest over the last few years. Recent studies in VQA research can be roughly divided into the following categories: better visual features [25, 26, 27], more powerful model architectures [28, 29, 30, 18], and more effective learning paradigms [31, 32, 33, 34, 35]. Most current state-of-the-art VQA methods employ the Transformer architecture [36]. By incorporating vision-language pretraining on large-scale datasets, they have approached or even surpassed human-level performance on several representative benchmarks [2, 3, 4, 37, 38]. Besides these studies on general-purpose VQA, there is also a growing trend towards exploring more granular VQA tasks with specific reasoning skills, e.g., neural-symbolic reasoning [39, 40] and knowledge utilization [5, 7].
Knowledge-based VQA. The core of this task lies in knowledge acquisition and integration. Early explorations parse the inputs into structured queries and retrieve supporting knowledge from fixed knowledge bases (KBs) to obtain the answers [5, 6]. As the provided knowledge resources are not sufficient to represent general knowledge, subsequent research mainly focuses on acquiring explicit knowledge from multiple open-domain knowledge resources, e.g., ConceptNet [9], Wikipedia [41], and Google Images [12]. This retrieved knowledge is integrated with the image-question pair for answer prediction [12, 42, 13]. Motivated by the powerful capacities of LLMs (e.g., GPT-3 [1]) in knowledge reasoning, recent state-of-the-art approaches regard an LLM as an implicit knowledge engine. They either utilize it to predict answers from given questions and extract visual captions [15] or to extract answer candidates with evidence to improve answer prediction [16, 17]. Nevertheless, they have not fully activated the reasoning capability of LLMs, as the necessary visual information to answer the question is not represented exactly. This motivates us to explore the strategies for prompting LLMs with question-aware information (i.e., answer heuristics). Similar to Prophet, a concurrent work PromptCap also aims to enhance the input information for LLMs by learning a question-aware captioning model [43]. However, PromptCap needs to use LLM in both the training and testing phases, which incurs tremendous computational costs as the training set is usually large. In contrast, Prophet is more economical as it only utilizes LLM in the testing phase.
@@
3 The Prophet Framework
[…]
@@ 3.1 Preliminaries
GPT-3 is an autoregressive language model pretrained on a tremendous dataset. During inference, in-context few-shot learning formulates a new downstream task as a text sequence generation task on the frozen model. Given a testing input $\boldsymbol{x}$, its target $\boldsymbol{y}$ is predicted conditioned on a formatted prompt $\boldsymbol{p}(\boldsymbol{h},\mathcal{E},\boldsymbol{x})$, where $\boldsymbol{h}$ refers to a prompt head, aka instruction, that describes the task, =$\mathcal{E}=\{\boldsymbol{e}_{1},\boldsymbol{e}_{2},...,\boldsymbol{e}_{n}\}$ corresponds to $n$ in-context examples. Let the target $\boldsymbol{y}=(y^{1},y^{2},...,y^{L})$ be a text sequence of $L$ tokens. For notational convenience, we denote $[l]$ as a set of natural numbers from 1 to $l$ and use $\boldsymbol{y}^{[l]}=(y^{1},...,y^{l})$ to represent a sub-sequence containing the first $l$ words of $\boldsymbol{y}$. At each decoding step $l$, we have:
@@ 3.2 Stage-1: Answer Heuristics Generation
As shown in Fig. 3, existing VQA methods can be categorized into discriminative and generative ones based on the ways they obtain answers. This discrepancy leads to different strategies for answer heuristics generation. We elaborate the strategy for each of the two classes of VQA models below.
Figure 3: Discriminative vs. generative VQA models. Taking an image (V) and a question (Q) as inputs, a typical discriminative VQA model like MCAN [18] performs multi-class classification to predict the most relevant answer (may contain multiple words) from a predefined answer vocabulary, while a typical generative VQA model like mPLUG [19] iteratively predicts one answer word at a time to constitute the final answer.
@@ 3.2.1 Discriminative VQA models
Answer-aware examples. Several previous studies have shown that the choice of in-context examples is crucial for GPT-3’s few-shot learning performance [15]. Their results motivate us to devise an answer-aware example selection strategy.
@@ 3.2.2 Generative VQA models
where each $\boldsymbol{w}_{j}$ represents an answer consisting of a sequence of answer words and $s_{j}\in\mathbb{R}^{+}$ denotes its corresponding confidence score calculated over all the answer words. The answer candidate set $\mathcal{C}$ is obtained from the generative model $\mathcal{M}_{\mathrm{gen}}$ equipped with the beam search strategy. Specifically, we initialize each answer $\boldsymbol{w}_{j}$ with the same $[\mathtt{BOS}]$ token. At each decoding step $l$, each $\boldsymbol{w}_{j}$ of length $l$ is first passed through $\mathcal{M}_{\mathrm{gen}}$ to obtain its top-$K$ candidate words with the highest scores. After that, an expand-then-reduce strategy is performed to update the $K$ answers: (i) expand step: each $\boldsymbol{w}_{j}$ is expanded $K$ times to combine with the $K$ candidate words, resulting in $K*K$ new candidates answers of length $l+1$; (ii) reduce step: among the $K*K$ candidate answers, only the top-$K$ ones with the highest accumulated scores $s=\sum_{i=1}^{l}\mathrm{log}~{}y^{i}$ are retained, which are then regarded as the inputs to the next decoding step.
@@
Using the score function above, we obtain the top-$N$ nearest neighbors of the query input in the training set and then format them as the answer-aware examples $\mathcal{E}$ as follows:
@@ 3.3 Stage-2: Heuristics-enhanced Prompting
[…]

### v4

@@ Abstract
Knowledge-based visual question answering (VQA) requires external knowledge beyond the image to answer the question. Early studies retrieve required knowledge from explicit knowledge bases (KBs), which often introduces irrelevant information to the question, hence restricting the performance of their models. Recent works have resorted to using a powerful large language model (LLM) as an implicit knowledge engine to acquire the necessary knowledge for answering. Despite the encouraging results achieved by these methods, we argue that they have not fully activated the capacity of the blind LLM as the provided textual input is insufficient to depict the required visual information to answer the question. In this paper, we present Prophet—a conceptually simple, flexible, and general framework designed to prompt LLM with answer heuristics for knowledge-based VQA. Specifically, we first train a vanilla VQA model on a specific knowledge-based VQA dataset without external knowledge. After that, we extract two types of complementary answer heuristics from the VQA model: answer candidates and answer-aware examples. The two types of answer heuristics are jointly encoded into a formatted prompt to facilitate the LLM’s understanding of both the image and question, thus generating a more accurate answer. By incorporating the state-of-the-art LLM GPT-3 [1], Prophet significantly outperforms existing state-of-the-art methods on four challenging knowledge-based VQA datasets. Prophet is general that can be instantiated with the combinations of different VQA models (i.e., both discriminative and generative ones) and different LLMs (i.e., both commercial and open-source ones). Moreover, Prophet can also be integrated with modern large multimodal models in different stages, which is named Prophet++, to further improve the capabilities on knowledge-based VQA tasks.
@@
Visual question answering (VQA), knowledge-based VQA, large language models (LLMs), large multimodal models.
@@ 1 Introduction
Recent advances in multimodal learning have achieved remarkable progress in various vision-language tasks, including visual captioning [2, 3], visual grounding [4, 5, 6], and visual question answering (VQA) [7, 8]. Among these tasks, VQA poses unique challenges by requiring machines to answer free-form questions through reasoning about given images. Benefiting from large-scale vision-language pretraining [9, 10, 6], the state-of-the-art methods have even surpassed human level on several representative benchmarks [11, 12]. Despite the success of these methods, their reasoning abilities are far from satisfactory, especially when external knowledge is required to answer the questions. In this situation, the task of knowledge-based VQA is introduced to validate models’ abilities to leverage external knowledge. Early knowledge-based VQA benchmarks additionally provide structured knowledge bases (KBs) and annotate required knowledge facts for all the questions [13, 14]. More recently, benchmarks emphasizing on open-domain knowledge have been established [15, 16], which means KBs are no longer provided and any external knowledge resource can be used for answering. We focus on the task with open-domain knowledge in this paper.
A straightforward solution for knowledge-based VQA is to retrieve knowledge entries from explicit KBs, e.g., Wikipedia and ConceptNet [17]. Then, a KB-augmented VQA model performs joint reasoning over the retrieved knowledge, image, and question to predict the answer [18, 19, 20]. However, the performance of these retrieval-based approaches is limited for two reasons: (i) the required knowledge may not be successfully retrieved from the KBs; and (ii) even if the required knowledge is retrieved, plenty of irrelevant knowledge is inevitably introduced, which hampers the learning of VQA models.
@@
Figure 1: Conceptual comparisons of three knowledge-based VQA frameworks using a frozen LLM model, e.g., GPT-3 [1]. While PICa [21], KAT [22], and REVIVE [23] directly feed the caption (C) and question (Q) into the LLM as the prompt, we argue that the information they provide for the LLM is insufficient thus cannot fully activate the LLM’s potential. In contrast, our Prophet learns a vanilla VQA model without external knowledge to produce answer heuristics, which endows the LLM with richer and more task-specific information for answer prediction. In contrast to the counterparts that resort to specific VQA models and LLMs, Prophet is general that can be instantiated with the combinations of different VQA models (i.e., both discriminative and generative ones) and different LLMs (i.e., both commercial and open-source ones). Moreover, Prophet can also be integrated with large multimodal models (LMMs) in different stages, which is termed Prophet++, to further improve the capabilities on knowledge-based VQA tasks.
@@
    The captions cannot cover all the necessary information in the image. Consider the example in Fig. 1, the caption “a group of people walk in a city square” contributes nothing to answering the question “what fruit comes from these trees”. In such situation, the LLM has to make an aimless and biased guess to answer the question.
@@
    LLMs like GPT-3 employ a few-shot learning paradigm that requires a few in-context examples to adapt to new tasks. Therefore, the choice of these examples is critical to model performance. As reported in [21], existing example selection strategies achieve far inferior performance to the oracle strategy that selects examples for a testing sample based on the similarity of their ground-truth answers, which is unavailable during testing.
@@
In this paper, we present Prophet—a conceptually simple yet effective framework designed to prompt LLMs with answer heuristics for knowledge-based VQA. By answer heuristics, we mean some promising answers that are presented in a proper manner in the prompt. Specifically, we introduce two types of complementary answer heuristics, namely answer candidates and answer-aware examples, to overcome the limitations in (i) and (ii), respectively. Given a testing input consisting of an image and a question, the answer candidates refer to a list of promising answers to the testing input, where each answer is associated with a confidence score. The answer-aware examples refer to a list of in-context examples, where each example has a similar answer to the testing input. Fortunately, these two types of answer heuristics can be simultaneously obtained from any vanilla VQA model trained on a specific knowledge-based VQA dataset. A schematic of Prophet is illustrated in Fig. 1.
@@
A preliminary version of this manuscript was published in [24]. Based on that version, we have made the following contributions to further improve the capabilties and generality of Prophet: (i) we investigate diverse types of VQA models, including the classical discriminative models trained from scratch and the latest generative VQA models pretrained on large-scale corpus; (ii) we expand the used LLM from the commercial GPT-3 model to a wide range of open-source models; (iii) apart from OK-VQA and A-OKVQA, we conduct more experiments on two other knowledge-based VQA datasets, namely ScienceQA [25] and TextVQA [26]. Furthermore, we present an improved Prophet++ framework by introducing large multimodal models (LMMs) into Prophet, which generates a new type of answer heuristic (i.e., answer-aware rationales) in the first stage and enables an extra visual input in the second stage.
The source code is made available here [^1] https://github.com/MILVLG/prophet. We hope our studies may inspire future research on knowledge-based VQA and universal vision-language learning.
@@ 2 Related Work
Visual Question Answering (VQA). VQA has been of growing interest over the last few years. Recent studies in VQA research can be roughly divided into the following categories: better visual features [2, 27, 28], more powerful model architectures [7, 29, 8], and more effective learning paradigms [30, 31, 9, 32, 33]. Most current state-of-the-art VQA methods employ the Transformer architecture [34]. By incorporating vision-language pretraining on large-scale datasets, they have approached or even surpassed human-level performance on several representative benchmarks [11, 12]. Besides these studies on general-purpose VQA, there is also a growing trend towards exploring more granular VQA tasks with specific reasoning skills, e.g., neural-symbolic reasoning [35, 36] and knowledge utilization [13, 15].
Knowledge-based VQA. The core of this task lies in knowledge acquisition and integration. Early explorations parse the inputs into structured queries and retrieve supporting knowledge from fixed knowledge bases (KBs) to obtain the answers [13, 14]. As the provided knowledge resources are not sufficient to represent general knowledge, subsequent research mainly focuses on acquiring explicit knowledge from multiple open-domain knowledge resources, e.g., ConceptNet [17], Wikipedia [37], and Google Images [19]. This retrieved knowledge is integrated with the image-question pair for answer prediction [19, 20]. Motivated by the powerful capacities of LLMs (e.g., GPT-3 [1]) in knowledge reasoning, recent state-of-the-art approaches regard an LLM as an implicit knowledge engine. They either utilize it to predict answers from given questions and extract visual captions [21] or to extract answer candidates with evidence to improve answer prediction [22, 23]. Nevertheless, they have not fully activated the reasoning capability of LLMs, as the necessary visual information to answer the question is not represented exactly.
This motivates us to explore the strategies for prompting LLMs with question-aware information (i.e., answer heuristics). Similar to Prophet, a concurrent work PromptCap also aims to enhance the input information for LLMs by learning a question-aware captioning model [38]. However, PromptCap needs to use LLM in both the training and testing phases, which incurs tremendous computational costs as the training set is usually large. In contrast, Prophet is more economical as it only utilizes LLM in the testing phase.
@@
3 The Prophet and Prophet++ Frameworks
As depicted in Fig. 2, our Prophet is a conceptually simple two-stage framework. In the answer heuristics generation stage, a vanilla VQA model is learned to generate two types of answer heuristics, i.e., answer candidates and answer-aware examples (detailed in §$\S$3.2). In the heuristics-enhanced prompting stage, the answer heuristics, question, and caption are integrated into a formatted prompt to instruct a frozen LLM to predict an answer (detailed in §$\S$3.3). Moreover, we introduce modern large multimodal models (LMMs) in different stages of Prophet to obtain a more powerful Prophet++ framework, which is detailed in §$\S$3.4.
@@ 3.1 Preliminaries
GPT-3 is an autoregressive language model pretrained on a tremendous dataset. During inference, in-context few-shot learning formulates a new downstream task as a text sequence generation task on the frozen model. Given a testing input $\boldsymbol{x}$, its target $\boldsymbol{y}$ is predicted conditioned on a formatted prompt $\boldsymbol{p}(\boldsymbol{h},\mathcal{E},\boldsymbol{x})$, where $\boldsymbol{h}$ refers to a prompt head, aka instruction, that describes the task, =$\mathcal{E}=\{\boldsymbol{e}_{1},...,\boldsymbol{e}_{n}\}$ corresponds to $n$ in-context examples. Let the target $\boldsymbol{y}=(y^{1},...,y^{L})$ be a text sequence of $L$ tokens. For notational convenience, we denote $[l]$ as a set of natural numbers from 1 to $l$ and use $\boldsymbol{y}^{[l]}=(y^{1},...,y^{l})$ to represent a sub-sequence containing the first $l$ words of $\boldsymbol{y}$. At each decoding step $l$, we have:
@@ 3.2 Stage-1: Answer Heuristics Generation
As shown in Fig. 3, existing VQA methods can be categorized into discriminative and generative ones based on the ways they obtain answers. This discrepancy leads to different strategies for answer heuristics generation. We elaborate the strategy for each class of VQA models below.
Figure 3: Discriminative vs. generative VQA models. Taking an image (V) and a question (Q) as inputs, a typical discriminative VQA model performs multi-class classification to predict the most relevant answer (may contain multiple words) from a predefined answer vocabulary, while a typical generative VQA model iteratively predicts one answer word at a time to constitute the final answer.
@@ 3.2.1 Discriminative VQA models
Answer-aware examples. Several previous studies have shown that the choice of in-context examples is crucial for LLM’s few-shot learning [21]. Their results motivate us to devise an answer-aware example selection strategy.
@@ 3.2.2 Generative VQA models
where each $\boldsymbol{w}_{j}$ represents an answer consisting of a sequence of answer words and $s_{j}\in\mathbb{R}^{+}$ denotes its corresponding confidence score calculated over all the answer words. The answer candidate set $\mathcal{C}$ is obtained from the generative model $\mathcal{M}_{\mathrm{gen}}$ equipped with the beam search strategy.
Specifically, we initialize each answer $\boldsymbol{w}_{j}$ with the same $[\mathtt{BOS}]$ token. At each decoding step $l$, each $\boldsymbol{w}_{j}$ of length $l$ is first passed through $\mathcal{M}_{\mathrm{gen}}$ to obtain its top-$K$ candidate words with the highest scores. After that, an expand-then-reduce strategy is performed to update the $K$ answers: (i) expand step: each $\boldsymbol{w}_{j}$ is expanded $K$ times to combine with the $K$ candidate words, resulting in $K*K$ new candidates answers of length $l+1$; (ii) reduce step: among the $K*K$ candidate answers, only the top-$K$ ones with the highest accumulated scores $s=\sum_{i=1}^{l}\mathrm{log}~{}y^{i}$ are retained, which are then regarded as the inputs to the next decoding step.
@@
Using Eq. (12), we obtain the top-$N$ nearest neighbors of the query input in the training set and then format them as the answer-aware examples $\mathcal{E}$ as follows:
@@ 3.3 Stage-2: Heuristics-enhanced Prompting
Our in-context examples are derived from the obtained $N$ answer-aware examples =$\mathcal{E}=\{\boldsymbol{e}_{1},\boldsymbol{e}_{2},...,\boldsymbol{e}_{N}\}$. Based on PICa’s template in §$\S$3.1, for example $\boldsymbol{e}_{i}$, we introduce its answer candidates $\mathcal{C}_{i}$ by adding one line of code as follows:
@@
3.4 Prophet++ with Frozen Large Multimodal Models
The original design of Prophet is based on pure-text LLMs that cannot perceive the images. Recently, a series of large multimodal models (LMMs), e.g., GPT-4V, GPT-4o, and LLaVA [42], have been proposed and show remarkable capabilities on various multimodal tasks. This raises a natural question: “Can these powerful LMMs be utilized in Prophet to further facilitate its capabilities?” To this end, we propose Prophet++, an extended framework that utilizes LMMs in both stages of Prophet, as shown in Fig 4.
In the first stage of Prophet++, we first use the VQA model to extract the two aforementioned answer heuristics in Prophet. Inspired by the strong reasoning capability of the latest LMMs, we extract a new type of answer heuristic, i.e., answer-aware rationales, which contains the reasoning process to answer the question. The prompt to generate answer-aware rationales is as follows:
You are provided with an image and a question. Please
think step by step to generate the rationale to help
answering the question.
where a few examples are also provided to control the output format. The output rationale is appended to the extracted image caption to be used in the second stage.
In the second stage of Prophet++, we replace the frozen LLM in Prophet with an LMM to handle multimodal inputs. Note that it takes a considerable amount of tokens to encode an image by a typical LMM. To restrict the total length of the multimodal inputs, only the test image is fed into the LMM while the images of the in-context examples are omitted.
@@ 4 Experiments
Figure 4: The Prophet++ framework additionally introduces large multimodal models (LMMs) in different stages of Prophet. Specifically, the LMM in stage-1 is used to generate a new type of answer heuristic, i.e., the answer-aware rationales while the LMM in stage-2 is used to handle an extra visual input (V), thus providing more comprehensive knowledge to answer the question.